OpenStudy (anonymous):

Solve the following word problem by using an equation and factoring. The sum of a positive number and its square is 20. What is the number? A. 1 B. 2 C. 3 D. 4 E. 5

OpenStudy (anonymous):

x+x^2=20 x^2+x-20=0 x^2+5x-4x-20=0 x(x+5)-4(x+5)=0 (x+5)(x-4)=0 two solutions: x+5=0 or x-4=0 Since x is positve, consider second solution giving x=4

OpenStudy (anonymous):

what ^^ said! :)